# API Pack: A Massive Multi-Programming Language Dataset for API Call Generation

### v5

@@ 1 Introduction
Large language models (LLMs) have shown promise in assisting with software engineering tasks (Li et al., 2023a; Ozkaya, 2023; Fan et al., 2023; Belzner et al., 2023; Ross et al., 2023; Hou et al., 2023; Ebert & Louridas, 2023; Chang et al., 2024; Wang et al., 2024), with a primary focus on code generation (Chen et al., 2021b; Xu et al., 2022; Sarsa et al., 2022; Vaithilingam et al., 2022; Poesia et al., 2022; Wang et al., 2023b; Shrivastava et al., 2023a; Liang et al., 2023; Zan et al., 2023; Liu et al., 2023a; Shrivastava et al., 2023b; Wei et al., 2023; Muennighoff et al., 2023; Li et al., 2023c; Huang et al., 2023; Lozhkov et al., 2024; Thakur et al., 2024; Mishra et al., 2024). While these advances are significant, developers still face a time-consuming challenge: finding and adapting API code examples from lengthy documentation (Meng et al., 2018). This manual process of searching for appropriate code examples remains cumbersome and inefficient. Although prior research has addressed API intent detection (finding the right API for a task), the more complex challenge of generating specific API call code has received limited attention, particularly given the varying syntax of HTTP protocol elements across programming languages.
To achieve this goal, we create API Pack, a dataset designed to improve the API call generation capabilities of LLMs. With over 1 million instances spanning 10 programming languages, API Pack represents the largest open-source instruction dataset for API call generation and intent detection (see Table 1). Our dataset distinguishes itself from prior works (Zan et al., 2022; Zhang et al., 2023a; Xu et al., 2023b; Patil et al., 2023; Qin et al., 2023) in two key aspects: programming language diversity and dataset size. By including API calls across multiple languages, API Pack enables the first comprehensive study of cross-language skill transfer in API generation, i.e., examining how improvements in one language generalize to others. Additionally, its extensive collection of real-world API examples allows for the evaluation of generalization capabilities through controlled variations in training data volume.
@@
This paper is structured as follows: We begin with a review of related work (Section 2), followed by a detailed description of the API Pack dataset construction (Section 3). Our experimental method for model fine-tuning and evaluation is presented in Section 4, with key findings discussed in Section 5. We conclude in Section 6 with a discussion of limitations and future directions. Additional considerations regarding ethics and broader impacts are addressed in Appendices A.1 and A.2, respectively. The API Pack dataset is made available under the Creative Commons Attribution 4.0 International License, with accompanying source code released under the MIT License.
@@ 2.1 Methods to Generate Instruction Data with LLMs
The manual creation of instruction datasets is a time-consuming and resource-intensive process (Xu et al., 2023a). To address this challenge, researchers have developed automated methods leveraging LLMs. Two notable approaches in this domain are Self-Instruct (Wang et al., 2023a) and Evol-Instruct (Xu et al., 2023a), both of which focus on generating and filtering instruction-response pairs. Self-Instruct utilizes in-context learning to modify seed examples and employs ROUGE-L similarity and heuristics for filtering. In contrast, Evol-Instruct uses targeted prompting for data generation and relies on predefined heuristics for filtering.
For code-specific datasets, hybrid LLM-based pipelines offer an alternative approach. These pipelines typically combine high-quality code samples from human with LLM-generated instructions (Patil et al., 2023; Xu et al., 2023b). While human verification remains the gold standard Wang et al. (2023a), it becomes impractical at scale. Recent work has instead adopted powerful LLMs like GPT-4 and Mistral as automated judges, using example-guided heuristics derived from human-reviewed data (Liu et al., 2023c). Similar techniques have been applied to evaluate instruction complexity (Chen et al., 2023; Lu et al., 2023).
@@ 3 API Pack
API Pack is an instruction dataset with over one million instances. Each instance contains an input-output pair plus additional information about the API and respective endpoints. The inputs are instructions for finding an API call to solve a coding task, including a task description in software engineering language and the name of the API. The outputs are API call examples, specifically HTTP request code snippets curated from OpenAPI specification (OAS) files. The data in API Pack is sourced from four hubs that store OAS files: RapidAPI [^1] https://rapidapi.com/categories, APIGurus [^2] https://apis.guru/, Swaggerhub [^3] https://app.swaggerhub.com/search, and IBM’s public API Hub [^4] https://developer.ibm.com/apis/. To the best of our knowledge, these API Hubs were the only sources with publicly available OAS files at a large scale at the time we collected data. From these sources, we kept all the APIs that passed our filtering criteria (see Section 3.1). Table 2 summarizes the total number of APIs, unique endpoints, and total instances that are part of API Pack.
@@ 3.1 Data Pre-processing
This process involves two main steps. First, we filter out OAS files with non-English metadata (less than 1%) and those with zero endpoints, as they cannot be used to generate API calls. Next, we extract key information from the remaining OAS files. At the API level, we collect the name, description, and provider. At the endpoint level, we extract the name, functionality, description, method, and path. To ensure data quality, we remove instances missing critical information needed for generating API calls (e.g., method, path, or endpoint name) or instructions (e.g., functionality or description). This step ensures that all dataset instances include the necessary details for generating accurate API call examples and instructions.
@@ 3.3 Instructions Generation
This process involves creating high-quality instruction examples and generating instruction candidates based on them. First, we randomly select three endpoints from each API DB and use their details (e.g., functionality, description, endpoint name, path) along with the API name to fill predefined instruction templates. This results in three instruction examples per API DB file. For the API Gurus and IBM API Hub sources, three co-authors of this work manually refine the generated instruction examples. The refinement criteria are: 1) correcting grammatical errors, 2) removing unnecessary information, and 3) ensuring the API name is included and correct in the examples. Most annotations were done independently, but annotators discussed non-standard cases to reach consensus. To streamline the refinement process for Swaggerhub and RapidAPI sources, we use an LLM (Mistral-7B-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2) instead of human effort. We create a prompt (see Prompt 2 in Appendix A.11) based on manual review heuristics and use it to task the LLM with refining instruction examples for these sources. Next, we use the high-quality instruction examples to generate five instruction candidates for each API DB instance. We prompt the LLM with the endpoint’s information and provide the high-quality instruction examples as in-context examples (see Prompt 3 in Appendix A.11). Figure 7 in Appendix A.8 shows an instance with one instruction candidate and its corresponding API call. Appendix A.9 presents all five candidates generated for the same instance.
@@ A.1 Ethical and Considerations
  • Hallucination: API Pack’s code (API calls) was not generated via an LLM. Thus, there is no risk of code errors in the dataset due to LLMs hallucination. That said, hallucination is a general risk associated with the use of LLMs. Therefore, code generated with LLMs that were fine-tuned with API Pack or any other code dataset must be always verified by a developer.
  • Data Ownership and License Information: The OAS files used to build API Pack come from public and official API data hubs. API data owners published these files on those websites themselves. That said, if a legitimate API code owner requires their data to be removed from API Pack we will do so. Regarding API’s license, only few OAS files contained this information. We are committed to tracking down complete and accurate license information for all APIs included in API Pack in a future version of the dataset.
@@ A.2 Broader Impact
  • Enhanced Software Productivity: The presented advancements promise significant software development acceleration by automating routine coding tasks. This could enable faster iterations, reduced costs, and potentially higher-quality products. However, reliance on auto-generated code risks reduced developer accountability and control. The oversight of human eye to ensure accuracy is still required.
@@ A.8 API Pack Dataset Instance Example
Figure 7 shows a fragment of an API Pack data instance. As data is sourced from public OAS files, personally identifiable and sensitive information in API calls is represented by anonymous strings (e.g., REPLACE_BASIC_AUTH), which indicates the need to pass this value as an argument in real execution. In this example, we only show one instruction for clarity, we represent the place for the other 4 instructions generated for this instance with the symbol … .

### v6

@@ 1 Introduction
Large language models (LLMs) have shown promise in assisting in software engineering tasks (Li et al., 2023a; Ozkaya, 2023; Fan et al., 2023; Belzner et al., 2023; Ross et al., 2023; Hou et al., 2023; Ebert & Louridas, 2023; Chang et al., 2024; Wang et al., 2024), with a primary focus on code generation (Chen et al., 2021b; Xu et al., 2022; Sarsa et al., 2022; Vaithilingam et al., 2022; Poesia et al., 2022; Wang et al., 2023b; Shrivastava et al., 2023a; Liang et al., 2023; Zan et al., 2023; Liu et al., 2023a; Shrivastava et al., 2023b; Wei et al., 2023; Muennighoff et al., 2023; Li et al., 2023c; Huang et al., 2023; Lozhkov et al., 2024; Thakur et al., 2024; Mishra et al., 2024). While these advances are significant, developers still face a time-consuming challenge: manually finding and adapting API code examples from lengthy documentation (Meng et al., 2018). Although prior research has addressed API intent detection (finding the right API for a task), the more complex challenge of generating specific API call code has received limited attention, particularly given the varying syntax of HTTP protocol elements across programming languages.
To achieve this goal, we create API Pack, a dataset designed to improve the API call generation capabilities of LLMs. With over 1 million instances spanning 10 programming languages, API Pack represents the largest open-source instruction dataset for API call generation and intent detection (see Table 1). Our dataset distinguishes itself from prior work (Zan et al., 2022; Zhang et al., 2023a; Xu et al., 2023b; Patil et al., 2023; Qin et al., 2023) in two key aspects: programming language diversity and dataset size. By including API calls across multiple languages, API Pack enables the first comprehensive study of cross-language skill transfer in API generation, as it examines how improvements in one language generalize to others. Additionally, its extensive collection of real-world API examples allows for the evaluation of generalization capabilities through controlled variations in training data volume.
@@
This paper is structured as follows: We begin with a review of related work (Section 2), followed by a detailed description of the API Pack dataset construction (Section 3). Our experimental method for model fine-tuning and evaluation is presented in Section 4, with key findings discussed in Section 5. We conclude in Section 6 with a discussion of limitations and future directions. Additional considerations regarding ethics and broader impact are addressed in Appendices A.1 and A.2, respectively. The API Pack dataset is made available under the Creative Commons Attribution 4.0 International License, with accompanying source code released under the MIT License.
@@ 2.1 Methods to Generate Instruction Data with LLMs
The manual creation of instruction datasets is a time-consuming and resource-intensive process (Xu et al., 2023a). To address this challenge, researchers have developed automated methods leveraging LLMs. Two notable approaches in this domain are Self-Instruct (Wang et al., 2023a) and Evol-Instruct (Xu et al., 2023a), both focused on generating and filtering instruction-response pairs. Self-Instruct utilizes in-context learning to modify seed examples and employs ROUGE-L similarity and heuristics for filtering. In contrast, Evol-Instruct uses targeted prompting for data generation and relies on predefined heuristics for filtering.
For code-specific datasets, hybrid LLM-based pipelines offer an alternative approach. These pipelines typically combine high-quality human code samples with LLM-generated instructions (Patil et al., 2023; Xu et al., 2023b). While human verification remains the gold standard Wang et al. (2023a), it becomes impractical at scale. Recent work has instead adopted powerful LLMs like GPT-4 and Mistral as automated judges, using example-guided heuristics derived from human-reviewed data (Liu et al., 2023c). Similar techniques have been applied to evaluate instruction complexity (Chen et al., 2023; Lu et al., 2023).
@@ 3 API Pack
API Pack [^1] https://huggingface.co/datasets/apipack/API-Pack-Dataset is an instruction dataset with over one million instances. Each instance contains an input-output pair plus additional information about the API and respective endpoints. The inputs are instructions for finding an API call to solve a coding task, including a task description in software engineering language and the name of the API. The outputs are API call examples, specifically HTTP request code snippets curated from OpenAPI specification (OAS) files. The data in API Pack is sourced from four hubs that store OAS files: RapidAPI [^2] https://rapidapi.com/categories, APIGurus [^3] https://apis.guru/, Swaggerhub [^4] https://app.swaggerhub.com/search, and IBM’s public API Hub [^5] https://developer.ibm.com/apis/. To the best of our knowledge, these API Hubs were the only sources with publicly available OAS files at a large scale at the time we collected data. From these sources, we kept all the APIs that passed our filtering criteria (see Section 3.1). Table 2 summarizes the total number of APIs, unique endpoints, and total instances that are part of API Pack.
@@ 3.1 Data Pre-processing
This process involves two main steps. First, we filter out OAS files with non-English metadata (less than 1%) and those with zero endpoints, as they cannot be used to generate API calls. Next, we extract key information from the remaining OAS files. At the API level, we collect the name, description, and provider. At the endpoint level, we extract the name, functionality, description, method, and path. To ensure data quality, we remove instances missing critical information needed for generating API calls (e.g., method, path, endpoint name) or instructions (e.g., functionality, description). This step ensures that all dataset instances include the necessary details for generating accurate API call examples and instructions.
@@ 3.3 Instructions Generation
This process involves creating high-quality instruction examples and generating instruction candidates based on them. First, we randomly select three endpoints from each API DB and use their details (e.g., functionality, description, endpoint name, path) along with the API name to fill predefined instruction templates. This results in three instruction examples per API DB file. For the API Gurus and IBM API Hub sources, three co-authors of this work manually refine the generated instruction examples. The refinement criteria are: 1) correcting grammatical errors, 2) removing unnecessary information, and 3) ensuring the API name is included and correct in the examples. Most annotations were done independently, but annotators discussed non-standard cases to reach consensus. To streamline the refinement process for Swaggerhub and RapidAPI sources, we use an LLM (Mistral-7B-Instruct-v0.2777https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2) instead of human effort. We create a prompt (see Prompt 2 in Appendix A.11) based on manual review heuristics and use it to task the LLM with refining instruction examples for these sources. Next, we use the high-quality instruction examples to generate five instruction candidates for each API DB instance. We prompt the LLM with the endpoint’s information and provide the high-quality instruction examples as in-context examples (see Prompt 3 in Appendix A.11). Figure 7 in Appendix A.8 shows an instance with one instruction candidate and its corresponding API call. Appendix A.9 presents all five candidates generated for the same instance.
@@ A.1 Ethical and Considerations
  • Hallucination: API Pack’s code (API calls) was not generated via an LLM. Thus, there is a low risk of code errors in the dataset due to LLMs hallucination. That said, hallucination is a general risk associated with the use of LLMs. Therefore, code generated with LLMs that were fine-tuned with API Pack or any other code dataset must be always verified by a developer.
  • Data Ownership and License Information: The OAS files used to build API Pack come from public and official API data hubs. API data owners published these files on those websites themselves. If a legitimate API code owner requires their data to be removed from API Pack we will do so. Regarding API’s license, only few OAS files contained this information. We are committed to tracking down complete and accurate license information for all APIs included in API Pack in a future version of the dataset.
@@ A.2 Broader Impact
  • Enhanced Software Productivity: The presented advancements promise significant software development acceleration by automating routine coding tasks. This could enable faster iterations, reduced cost, and potentially increase the quality of products. However, reliance on auto-generated code risks reduced developer accountability and control. The oversight of human eye to ensure accuracy is still required.
@@ A.8 API Pack Dataset Instance Example
Figure 7 shows a fragment of an API Pack data instance. As data is sourced from public OAS files, personally identifiable and sensitive information in API calls is represented by anonymous strings (e.g., REPLACE_BASIC_AUTH), which indicates the need to pass this value as an argument in real execution. In this example, we only show one instruction candidate for clarity.